OpenStudy (anonymous):

Rationalizing denominators- please walk me through the steps to solve this problem--

OpenStudy (anonymous):

\[8\sqrt{2}/\sqrt{3}\]

OpenStudy (anonymous):

the slash line is equivalent to a fraction line

OpenStudy (anonymous):

multiply the fraction by - sqrt3/-sqrt3 that is \[\sqrt{3} / \sqrt{3}\] that gives you \[8 \sqrt{6} /3\]

OpenStudy (anonymous):

got it. now final one is 8 over 3+\[\sqrt{7}\]

OpenStudy (anonymous):

do I multiply 8 by 3 and sqrt 7? and if so, what do I keep as denom?

OpenStudy (anonymous):

no the final answer is 8 multiplied by square root of 6 the whole thing over 3 my answer is the final answer because you are multiplying by root 3 / root3 which is 1 that does not change the value of the fraction so the answer i typed is the final answer

OpenStudy (anonymous):

where did you get sqrt 7 from ?

OpenStudy (anonymous):

that is not the answr, that is part of the problem: the problem is 8 OVER 3 PLUS sqrt7. haha sorry about that!

OpenStudy (anonymous):

let me know what answer you came up with

OpenStudy (anonymous):

hello?